# Breaking the Performance Ceiling in Complex Reinforcement Learning requires Inference Strategies

### v1

@@ 3.2 One budget, many uses: inference-time search and adaptation
These methods store information about partial solutions using past attempts to preferentially search promising regions of the solution space without updating the pre-trained policy. Simulation guided beam search (sgbs(Choo et al., 2022) provides the best time to performance balance in the literature, outperforming Monte Carlo Tree Search (Coulom, 2006). Like most tree searches, sgbs has three steps: expansion, simulation, pruning. Expansion uses the policy to decide on the most promising next actions (i.e., $)a=\text{top-}K(\pi_{\theta}(\cdot|o))$ from the current node (partial solution). A simulated rollout of an episode is produced greedily using $\pi_{\theta}$ and the return is collected for each node. Pruning keeps only the best nodes found so far based on the return. Solely the expansion step needs to be adapted for Multi-agent SGBS. This is trivial when the explicit joint actions are accessible (de Witt et al., 2020; Yu et al., 2022), since we can still select the top ones (i.e., $\boldsymbol{a}=\text{top-}K(\boldsymbol{\pi}_{\boldsymbol{\theta}}(\cdot|\boldsymbol{o}))$). For methods using auto-regressive action selection (Mahjoub et al., 2025), having access to the top joint actions is intractable, hence we sample K times stochastically from the same node (i.e., $\boldsymbol{a}[1],...,\boldsymbol{a}[K]\sim\boldsymbol{\pi}_{\boldsymbol{\theta}}(\cdot|\boldsymbol{o})$).
@@
These methods keep updating policy parameters at inference time. Given a base policy $\pi_{\theta}$, online fine-tuning optimises $\theta$ using inference-time rollouts and policy gradient updates: $\theta^{\prime}=\theta+\alpha\nabla_{\theta}J(\pi_{\theta})$, where $J(\pi_{\theta})$ represents an adaptation objective. In line with (Bello et al., 2016), we keep maximising expected returns (Bello et al., 2016) over past attempts on the fixed instance (instead of over a training distribution). Multi-agent online fine-tuning re-trains $\boldsymbol{\pi}_{\boldsymbol{\theta}}$ on the new instance using the MARL algorithm that was used during pre-training (Bello et al., 2016; Mahjoub et al., 2025; de Witt et al., 2020; Yu et al., 2022).
@@
These methods pre-train a collection of diverse specialised policies which can be used to search for the most appropriate solution at inference-time. compass (Chalumeau et al., 2023b) encodes specialised policies in a continuous latent space $\mathcal{L}$ by augmenting a pre-trained policy to condition on both the observation and a latent vector sampled from $\mathcal{L}$ (i.e., $a\sim\pi_{\theta}(\,\cdot\mid o,~{}z),~{}~{}z\sim\mathcal{L}$): effectively creating a continuous collection of policies. compass achieves SOTA in single-agent RL for CO. To avoid having the latent space of Multi-agent COMPASS growing exponentially with the number of agents, we keep one latent space $\mathcal{L}$ for all agents (i.e., $\boldsymbol{a}\sim\boldsymbol{\pi}_{\boldsymbol{\theta}}(\,\cdot\mid\boldsymbol{o},~{}z),~{}~{}z\sim\mathcal{L}$). This allows for tractable training, and for efficient inference search with the covariance matrix adaptation evolution strategy (cma-es(Hansen and Ostermeier, 2001). Aside from being multi-agent, we keep the training and inference phases close to the original method described in Chalumeau et al. (2023b).
@@ 4 Experiments
Figure 3: Overview of our tasks and experimental study.
@@
We use three MARL approaches to obtain our base policies: Independent PPO (de Witt et al., 2020) (ippo) and Multi-Agent PPO (Yu et al., 2022) (mappo), which are widely used and well-known MARL methods, and the recent SOTA sequence modelling approach sable (Mahjoub et al., 2025). Each of these, referred to as base policies, is evaluated with all four inference strategies introduced in Section 3.2, namely stochastic sampling, sgbs, online fine-tuning and compass.
@@
Mahjoub et al. (2025) established SOTA over the most comprehensive MARL benchmark published in the field to date. Interestingly, their results demonstrate that there remain certain tasks for which no existing method (including sable) is able to achieve good performance. Specifically, these are tiny-2ag-hard, tiny-4ag-hard, small-4ag, small-4ag-hard, medium-4ag, medium-4ag-hard, medium-6ag, large-4ag, large-4ag-hard, large-8ag, large-8ag-hard, xlarge-4ag and xlarge-4ag-hard from Multi-Robot Warehouse (Papoudakis et al., 2021) (RWARE), smacv2_10_units and smacv2_20_units from the StarCraft Multi-Agent challenge (Samvelyan et al., 2019; Ellis et al., 2023) (SMAC), and con-10x10x10a and con-15x15x23a from Connector (Bonnet et al., 2023).
Figure 4: Training the SOTA algorithm Sable to convergence is not enough to achieve optimal zero-shot performance.
Each environment (illustrated on Fig. 3) introduces distinct challenges that contribute to its complexity. RWARE requires agents to coordinate in order to pick up and deliver packages without collision and has a very sparse reward signal. In SMACv2 tasks, a team cooperates in real-time combat against enemies across diverse scenarios with randomised generation. Connector models the routing of a printed circuit board where agents must connect to designated targets without crossing paths. All three environments feature combinatorial and high-dimensional action spaces, partial observability and the need for tightly coordinated behaviours, making these 17 tasks a compelling test-bed for complex RL with desirable properties modelling aspects of real-world tasks. We use JAX-based implementations of Connector and RWARE from Jumanji (Bonnet et al., 2023) and for SMAC from JaxMARL (Rutherford et al., 2023).
@@
To obtain clear performance ceilings for each algorithm and best isolate the effects of inference strategies, we train all base policies until convergence. For the sake of continuity with previous work with truncated training budgets, typically of $20$M steps, we report the zero-shot results for each converged checkpoint compared to its previous corresponding reported performance on Fig. 4. We observe that in most tasks (14 out of 17), the converged policy stays below 70% normalised performance. compass requires an additional training phase, which reincarnates the existing base policies to create the latent space specialisation. For each base policy and task, we also train the compass checkpoint until convergence. This leads to $102$ trained policy checkpoints.
@@
Evaluating inference strategies in a way that is unbiased and aligned with real-world settings is challenging. Most papers report results where the budget is based on a number of attempts, hence not directly incorporating the time cost of the strategy. The time costs are reported, but it is tough to analyse due to the plurality of hardware used to obtain them. Having re-implemented all of the baselines in the same code base and setting the budget in terms of time (in seconds), we can avoid this bias in our study. We use the same fixed hardware for all our experiments, namely a NVIDIA-A100-SXM4-80GB GPU. For statistical robustness, we always run $128$ independent seeds. In all cases, we control for $\text{B}_{\text{max}}$ by varying the permitted number of batched parallel attempts instead of altering hardware between experiments. For aggregation across multiple tasks we follow the recommendations made by Agarwal et al. (2021) and use the rliable library to compute and report the inter-quartile mean (IQM) and 95% stratified bootstrap confidence intervals.
@@
To train the base policies, we re-use the hyperparameters reported in Mahjoub et al. (2025), which have been optimised for our tasks. For the inference strategies, we follow recommendations from the literature (Choo et al., 2022; Chalumeau et al., 2023b). All hyperparameters choices are reported in Appendix C.
@@ 4.1 A couple of seconds is all you need
In this section, we demonstrate that inference-time search can help reach close to maximum task performance, using base policies for which zero-shot performance stagnates around $60$%.
@@
To demonstrate that inference-time strategies are accessible, we use a small budget: $30$ seconds, and a compute capacity enabling to generate $64$ solutions in parallel. Each base policy is evaluated greedily for a single attempt, and then evaluated with the search budget, using each inference strategy. We report the performance distribution over the $17$ tasks in Fig. 5, and the performance gains offered by the best inference-time search over the best zero-shot on Fig. 1.
@@ Discussion
We can draw four main conclusions. First, inference-time search does provide a massive performance boost over zero-shot, which stands for every base policy. For the SOTA zero-shot method sable, this translates to pushing the best-ever achieved aggregated performance by more than $45$% and creating a system (sable +compass) that achieves close to $100$% win-rate in all tasks where this metric is available. Second, the improvement enabled over zero-shot performance increases significantly (almost exponentially) with respect to the complexity of the task (see Fig. 1). This suggests substantial gains are still ahead as the field moves toward increasingly realistic scenarios. Third, we observe that compass is the leading strategy across tasks and base algorithms, and that sable remains the SOTA base policy even when using inference-time search. Interestingly, under a small time budget, stochastic sampling outperforms tree search and online fine-tuning. We nevertheless show in the following section that this order can be reversed when more budget is provided.
@@ 4.2 Mapping performance with compute and time budget
We choose a maximum time of $300$ seconds, and evaluate all inference strategies using the leading base policy (sable), with a compute budget of $\{4,8,16,32,64,128,256\}$. All in all, we have $4$ inference strategies, $7$ possible compute budgets, $17$ tasks, and $128$ seeds per task, leading to 60 9286092860\,92860 928 evaluated episodes. This constitutes the largest study released on inference strategies. We report these results using contour plots, where the $x$-axis is time, $y$-axis the number of parallel attempts allowed (our proxy for compute) and colour corresponds to the performance achieved (win-rate when accessible or min-max normalised return) going from dark purple (min) to yellow (max). We keep the $8$ hardest tasks, the lower half based on the zero-shot performance of the converged sable checkpoints (see Fig. 4), on Fig. 6 and defer remaining tasks to Appendix A.
@@ Discussion
As a sanity check, we remark that performance always increases (colours become lighter) when time or compute increases (going towards the upper right corner). We now highlight three main observations. First, compass demonstrates impressive versatility and achieves significant gains over other inference strategies, dominating all maps, except for con-10x10x10a, where it gets slightly outperformed by online fine-tuning. Second, we observe high variance for online fine-tuning: getting close to compass for large budgets on con-10x10x10a, yet struggling to match stochastic sampling on others (e.g., RWARE’s large-8ag). This shows that fine-tuning can be detrimental by reducing the number of attempts made within the time budget. Plus, policy gradients can be unstable (small batch size) or converge to local optima. This observation disproves the common belief that inference-time search is as trivial as over-fitting to the problem instance. Finally, we observe that sgbs has good performance for high compute capacity and low time, but the greediness of its simulation step impacts its exploration capacity, preventing to benefit from additional time budget.
@@ 4.3 Scaling with increasing budget
In practical applications, time is often more restricted than compute: a couple of seconds or minutes can be allowed, sometimes a few hours (train scheduling for the next day), but rarely more. Being able to improve solutions’ quality under a fixed time budget by scaling compute is a valuable property. Consequently, methods that can scale with the compute available are particularly valuable. We look at this property in this section.
@@
We keep the time budget fixed, $300$ seconds, and we plot the final performance for each possible compute budget (still using the number of parallel attempts as a proxy). We use sable as the base policy and evaluate all the inference strategies across the $8$ hardest tasks. We report results per strategy, and aggregated over the tasks on Fig. 7.
Figure 7: Scaling of performance with respect to compute capacity, for a fixed time budget. Compute capacity ($x$-axis) is set by limiting the number of parallel attempts. sable +compass demonstrates impressive scaling properties, reaching the performance bounds of the benchmark.
@@ Discussion
As expected, stochastic sampling has the lowest scaling coefficient: its search solely relies on chance, no adaptation or additional search is happening. On the other hand, we can see that online fine-tuning benefits from more compute, probably due to a better estimation of the policy gradient. It nevertheless requires a budget of 64 parallel attempts to clearly outperform stochastic sampling. On the other hand, compass consistently provides a significant advantage. Its scaling trend seems linear at first, with a high coefficient, and only seems to decline for higher budgets, as performance limits are reached (over $95$% win-rate when accessible, best-ever observed performance elsewhere). We have a two-fold explanation for these particularly impressive scaling properties: (i) the diversity contained in the latent space can be exploited by more parallelism, leading to a massive exploration of the solution space, even when the initial policy is far from optimal, and (ii) the higher the batch size, the better the cma-es search, enabling compass to exploit even more information from any additional searching step allowed within the given time budget.
@@ 5 Conclusion
In this work, we demonstrate that inference-time strategies are a critical and underutilized lever for boosting the performance of RL systems in complex tasks, using multi-agent RL as a representative test-bed. While training-time improvements have long dominated the field, our results show that inference-time search may offer significant performance gains using only a few seconds to minutes of additional wall-clock time during execution. We introduce a unified framework for inference strategies, extend it to the multi-agent setting, and empirically validate its effectiveness under varying compute and time budgets. Our large-scale evaluation, the most comprehensive to date on inference-time methods, reveals three key takeaways: (i) inference-time search with a relevant strategy yields significant improvements, even under tight time constraints; (ii) the gains depend on the inference budget, and our contour maps provide practitioners with practical guidance based on their constraints; and (iii) sable +compass not only dominates the benchmark but also exhibits the most favourable scaling trends, making it particularly effective for increasingly complex decision-making problems.
@@
We focus on multi-agent RL to represent complex RL. We are aware that single-agent tasks can be arbitrarily complex. We believe that our main trends should generalise to the single-agent case, but leave it for future work. We intend two main future research directions. First, studying how to best combine existing paradigms together. Second, investigating how inference strategies reacts to evaluation out-of-distribution.

### v3

@@ 3.2 One budget, many possibilities: inference-time search and adaptation
These methods store information about partial solutions using past attempts to preferentially search promising regions of the solution space without updating the pre-trained policy. Simulation guided beam search (sgbs(choo2022simulationguided) provides the best time to performance balance in the literature, outperforming Monte Carlo Tree Search (coulom2006mcts). Like most tree searches, sgbs has three steps: expansion, simulation, pruning. Expansion uses the policy to decide on the most promising next actions (i.e., $a=\text{top-}K(\pi_{\theta}(\cdot|o))$ ) from the current node (partial solution). A simulated rollout of an episode is produced greedily using $\pi_{\theta}$ and the return is collected for each node. Pruning keeps only the best nodes found so far based on the return. Solely the expansion step needs to be adapted for Multi-agent SGBS. This is trivial when the explicit joint actions are accessible (dewitt2020independentlearningneedstarcraft; yu2022mappo), since we can still select the top ones (i.e., $\boldsymbol{a}=\text{top-}K(\boldsymbol{\pi}_{\boldsymbol{\theta}}(\cdot|\boldsymbol{o}))$). For methods using auto-regressive action selection (mahjoub2025sable), having access to the top joint actions is intractable, hence we sample K times stochastically from the same node (i.e., $\boldsymbol{a}[1],...,\boldsymbol{a}[K]\sim\boldsymbol{\pi}_{\boldsymbol{\theta}}(\cdot|\boldsymbol{o})$).
@@
These methods keep updating policy parameters at inference time. Given a base policy $\pi_{\theta}$, online fine-tuning optimises $\theta$ using inference-time rollouts and policy gradient updates: $\theta^{\prime}=\theta+\alpha\nabla_{\theta}J(\pi_{\theta})$, where $J(\pi_{\theta})$ represents an adaptation objective. In line with (Bello16), we keep maximising expected returns (Bello16) over past attempts on the fixed instance (instead of over a training distribution). Multi-agent online fine-tuning re-trains $\boldsymbol{\pi}_{\boldsymbol{\theta}}$ on the new instance using the MARL algorithm that was used during pre-training (Bello16; mahjoub2025sable; dewitt2020independentlearningneedstarcraft; yu2022mappo).
@@
These methods pre-train a collection of diverse specialised policies which can be used to search for the most appropriate solution at inference-time. compass (chalumeau2023combinatorial) encodes specialised policies in a continuous latent space $\mathcal{L}$ by augmenting a pre-trained policy to condition on both the observation and a latent vector sampled from $\mathcal{L}$ (i.e., $a\sim\pi_{\theta}(\,\cdot\mid o,~z),~~z\sim\mathcal{L}$): effectively creating a continuous collection of policies. compass achieves SOTA in single-agent RL for CO. To avoid having the latent space of Multi-agent COMPASS growing exponentially with the number of agents, we keep one latent space $\mathcal{L}$ for all agents (i.e., $\boldsymbol{a}\sim\boldsymbol{\pi}_{\boldsymbol{\theta}}(\,\cdot\mid\boldsymbol{o},~z),~~z\sim\mathcal{L}$). This allows for tractable training, and for efficient inference search with the covariance matrix adaptation evolution strategy cma-es (hansen2001cmaes). Aside from being multi-agent, we keep the training and inference phases close to the original method described in chalumeau2023combinatorial, and provide further details and explanations in Section˜F.2.
Unlike other inference strategies, compass includes an additional training phase, which remains accessible since the training phase is unconstrained (all policies are trained until convergence). Creating a COMPASS checkpoint from a pre-trained base policy involves adding parameters to process the latent vectors (Section˜F.2), resulting in a modest increase in model size. This increase never exceeds 22% of the total policy size, and has negligible impact on the overall computational or memory footprint. We report all parameter counts in Appendix˜I.
@@ 4 Experiments
Figure 3: Overview of our evaluation tasks and experimental study.
@@
We use three MARL approaches to obtain our base policies: Independent PPO (dewitt2020independentlearningneedstarcraft) (ippo) and Multi-Agent PPO (yu2022mappo) (mappo), which are widely used and well-known MARL methods, and the recent SOTA sequence modelling approach sable (mahjoub2025sable). Each of these, referred to as base policies, is evaluated with all four inference strategies introduced in Section˜3.2, namely stochastic sampling, sgbs, online fine-tuning and compass.
@@
mahjoub2025sable established SOTA over the most comprehensive MARL benchmark published in the field to date. Interestingly, their results demonstrate that there remain certain tasks for which no existing method (including sable) is able to achieve good performance. Specifically, these are tiny-2ag-hard, tiny-4ag-hard, small-4ag, small-4ag-hard, medium-4ag, medium-4ag-hard, medium-6ag, large-4ag, large-4ag-hard, large-8ag, large-8ag-hard, xlarge-4ag and xlarge-4ag-hard from Multi-Robot Warehouse (papoudakis2021benchmarking) (RWARE), smacv2_10_units and smacv2_20_units from the StarCraft Multi-Agent challenge (smac; ellis2023smacv2) (SMAC), and con-10x10x10a and con-15x15x23a from Connector (jumanji2023github).
Figure 4: Training the SOTA algorithm Sable to convergence is not enough to achieve optimal zero-shot performance. We report the mean normalised performance per task and report 95% bootstrap confidence intervals in Appendix˜L.
Each environment (illustrated on Fig.˜3) introduces distinct challenges that contribute to its complexity. RWARE requires agents to coordinate in order to pick up and deliver packages without collision and has a very sparse reward signal. In SMACv2 tasks, a team cooperates in real-time combat against enemies across diverse scenarios with randomised generation. Connector models the routing of a printed circuit board where agents must connect to designated targets without crossing paths. All three environments feature combinatorial and high-dimensional action spaces, partial observability and the need for tightly coordinated behaviours, making these 17 tasks a compelling test-bed for complex RL with desirable properties modelling aspects of real-world tasks. We use JAX-based implementations of Connector and RWARE from Jumanji (jumanji2023github) and for SMAC from JaxMARL (rutherford2023jaxmarl).
@@
To obtain clear performance ceilings for each algorithm and best isolate the effects of inference strategies, we train all base policies until convergence. For the sake of continuity with previous work with truncated training budgets, typically of 2020M steps, we report the zero-shot results for each converged checkpoint compared to its previous corresponding reported performance on Fig.˜4. We observe that in most tasks (14 out of 17), the converged policy stays below 70% normalised performance, demonstrating that the benchmark is still far from saturated. compass requires an additional training phase, which reincarnates the existing base policies to create the latent space specialisation. For each base policy and task, we also train the compass checkpoint until convergence. This leads to $102$ trained policy checkpoints.
@@
Evaluating inference strategies in a way that is unbiased and aligned with real-world settings is challenging. Most papers report results where the budget is based on a number of attempts, hence not directly incorporating the time cost of the inference strategy. The time costs are reported, but it is tough to analyse due to the plurality of hardware used to obtain them. Having re-implemented all of the baselines in the same code base and setting the budget in terms of time (in seconds), we can avoid this bias in our study. We use the same fixed hardware for all our experiments, namely a NVIDIA-A100-SXM4-80GB GPU. For statistical robustness, we always run $128$ independent seeds. In all cases, we control for $\text{B}_{\text{max}}$ by varying the permitted number of batched parallel attempts instead of altering hardware between experiments. For aggregation across multiple tasks we follow the recommendations made by agarwal2021rliable and use the rliable library to compute and report the inter-quartile mean (IQM) and 95% stratified bootstrap confidence intervals.
@@
To train the base policies, we re-use the hyperparameters reported in mahjoub2025sable, which have been optimised for our tasks. For the inference strategies, we follow recommendations from the literature (choo2022simulationguided; chalumeau2023combinatorial). All hyperparameters choices are reported in Appendix˜D.
@@ 4.1 A couple of seconds is all you need
In this section, we demonstrate that inference-time search can help reach close to maximum task performance, using base policies for which zero-shot performance stagnates around 6060%.
@@
To demonstrate that inference-time strategies are accessible, we use a small budget: $30$ seconds, and a compute capacity enabling to generate $64$ solutions in parallel. Each base policy is evaluated greedily for a single attempt, and then evaluated with the search budget, using each inference strategy. We report the performance distribution over the $17$ tasks in Fig.˜5, and the performance gains offered by the best inference-time search over the best zero-shot on Fig.˜1.
@@ Discussion
We can draw four main conclusions. First, inference-time search does provide a massive performance boost over zero-shot, which stands for every base policy. For the SOTA zero-shot method sable, this translates to pushing the best-ever achieved aggregated performance by more than 4545% and creating a system (sable +compass) that achieves close to 100100% win-rate in all tasks where this metric is available. Second, the improvement enabled over zero-shot performance increases significantly (almost exponentially) with respect to the complexity of the task (see Fig.˜1). This suggests substantial gains are still ahead as the field moves toward increasingly realistic scenarios. Third, we observe that compass is the leading strategy across tasks and base algorithms, and that sable remains the SOTA base policy even when using inference-time search. Interestingly, under a small time budget, stochastic sampling outperforms online fine-tuning. We nevertheless show in the following section that, given more budget, this result can be nuanced, and we share our interpretation of these findings.
@@ 4.2 Mapping performance with compute and time budget
We choose a maximum time of $300$ seconds, and evaluate all inference strategies using the leading base policy (sable), with a compute budget of $\{4,8,16,32,64,128,256\}$. All in all, we have $4$ inference strategies, $7$ possible compute budgets, $17$ tasks, and $128$ seeds per task, leading to 60 92860\,928 evaluated episodes. This constitutes the largest study released on inference strategies. We report these results using contour plots, where the $x$-axis is time, $y$-axis the number of parallel attempts allowed (our proxy for compute) and colour corresponds to the performance achieved (win-rate when accessible or min-max normalised return) going from dark purple (min) to yellow (max). We keep the $8$ hardest tasks, the lower half based on the zero-shot performance of the converged sable checkpoints (see Fig.˜4), on Fig.˜6 and defer remaining tasks to Appendix˜A.
@@ Discussion
As a sanity check, we remark that performance always increases (colours become lighter) when time or compute increases (going towards the upper right corner). We now highlight three main observations. First, compass demonstrates impressive versatility and achieves significant gains over other inference strategies, dominating all maps, except for con-10x10x10a, where it gets slightly outperformed by online fine-tuning. Second, we observe high variance for online fine-tuning: getting close to compass for large budgets on con-10x10x10a, yet struggling to match stochastic sampling on others (e.g., RWARE’s large-8ag). This shows that fine-tuning can be detrimental by reducing the number of attempts made within the time budget. Plus, policy gradients can be unstable (small batch size) or converge to local optima. This observation disproves the common belief that inference-time search is as trivial as over-fitting to the problem instance. Finally, we observe that sgbs, despite failing on the connector tasks, achieves competitive performance overall and even leads in the low-budget regime (i.e., below 100 seconds and fewer than 10 parallel attempts) in some scenarios (large-4ag and large-4ag-hard).
@@ 4.3 Scaling with increasing budget
In practical applications, time is often more restricted than compute: a couple of seconds or minutes can be allowed, sometimes a few hours (train scheduling for the next day), but rarely more. Being able to improve solutions’ quality under a fixed time budget by increasing compute is desirable. In this section, we analyse how different methods scale with additional compute.
Figure 7: Scaling of performance with respect to compute capacity (fixed time budget). Compute capacity is set by limiting the number of parallel attempts. sable +compass demonstrates impressive scaling properties, reaching the performance bounds.
@@
We keep the time budget fixed, $300$ seconds, and we plot the final performance for each possible compute budget (still using the number of parallel attempts as a proxy). We use sable as the base policy and evaluate all the inference strategies across the $8$ hardest tasks. We report results per strategy, and aggregated over the tasks on Fig.˜7.
@@ Discussion
As expected, stochastic sampling has the lowest scaling coefficient: its search solely relies on chance, no adaptation or additional search is happening. On the other hand, we can see that online fine-tuning benefits from more compute, probably due to a better estimation of the policy gradient. It nevertheless requires a budget of 64 parallel attempts to clearly outperform stochastic sampling. On the other hand, compass consistently provides a significant advantage. Its scaling trend seems linear at first, with a high coefficient, and only seems to decline for higher budgets, as performance limits are reached (over 9595% win-rate when accessible, best-ever observed performance elsewhere). We have a two-fold explanation for these particularly impressive scaling properties: (i) the diversity contained in the latent space can be exploited by more parallelism, leading to a massive exploration of the solution space, even when the initial policy is far from optimal, and (ii) the higher the batch size, the better the cma-es search, enabling compass to exploit even more information from any additional searching step allowed within the given time budget. We discuss this further in Appendix˜K.
@@ 5 Conclusion
In this work, we demonstrate that inference-time strategies are a critical and underutilized lever for boosting the performance of RL systems in complex tasks, using multi-agent RL as a representative test-bed. While training-time improvements have long dominated the field, our results show that inference-time search may offer significant performance gains using only a few seconds to minutes of additional wall-clock time during execution. We introduce a unified view of inference strategies, extend it to the multi-agent setting, and empirically validate its effectiveness under varying compute and time budgets.
Our large-scale evaluation, the most comprehensive to date on inference-time methods, reveals three key takeaways: (i) inference-time search with a relevant strategy yields significant improvements, even under tight time constraints; (ii) the gains depend on the inference budget, and our contour maps provide practitioners with practical guidance based on their constraints; and (iii) sable +compass not only dominates the benchmark but also exhibits the most favourable scaling trends, making it particularly effective for increasingly complex decision-making problems.
@@
We focus on multi-agent RL as it better captures the complexity of real-world decision-making systems, where successful operation often relies on coordination among multiple agents. Nonetheless, we acknowledge that naturally single-agent tasks can also be highly complex. In the spirit of seeding a wider investigation for future work beyond the multi-agent setting, we provide additional results on the single-agent Craftax benchmark in Appendix˜C, showing that a simple $30$-second inference-time search using stochastic sampling can provide a 3737% performance boost. This provides initial (but limited) evidence that our claims hold more generally and we leave a more thorough investigation for future work. Beyond the broader RL setting, we intend to investigate two main future research directions. First, studying how to best combine existing inference paradigms to leverage their complementary strengths. Second, investigating how inference strategies compare when evaluated out-of-distribution.